OpenStudy (anonymous):

Find an equation for the tangent line to the graph given by: -x^2-3xy+y^2=23 at the point (-2,3) please show the steps i wanna learn..

OpenStudy (anonymous):

Steps are find the implicit differentiation and then plug in x and y \[-2x-3y-3x(dy/dx)+2y(dy/dx)=(2x+3y)/(2y-3x) =5/12\]

OpenStudy (anonymous):

what happened to the 23 tho?

OpenStudy (anonymous):

and its asking to find the equation so what would the equation be?